OpenStudy (anonymous):

A scatter plot contains a trend line with a slope of -1/10. Does this trend line illustrate a correlation?

OpenStudy (anonymous):

A. Yes, there is a positive correlation. B. Yes, there is a negative correlation. C. No, there is no correlation. D. There is not enough information.

jimthompson5910 (jim_thompson5910):

is that all you're given? or are you given the correlation coefficient as well?

OpenStudy (anonymous):

thats all thats given

jimthompson5910 (jim_thompson5910):

ok well ideally you would have the correlation coefficient (or the data values to find the correlation coefficient) because this will help you determine how strong the linear correlation is

jimthompson5910 (jim_thompson5910):

but because this critical piece of info is missing, it's hard to say how strong the correlation is which leads me to think it's either B or D if the correlation is strong enough, then you'll have a negative correlation however, we don't have enough info (as I said above), which is why D could be an answer as well

jimthompson5910 (jim_thompson5910):

see what I'm saying?

OpenStudy (anonymous):

yea

jimthompson5910 (jim_thompson5910):

If I were answering this, I guess I would lean more towards D since that missing piece of info is very critical However, they could have easily implied that the correlation coefficient is assumed to be -0.5 or smaller, not sure though

OpenStudy (anonymous):

A scatter plot contains the points (-3, 7) and (-1, 2). Does this scatter plot illustrate a correlation? would this be a negative correlation?

jimthompson5910 (jim_thompson5910):

it only contains those points? and nothing else?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok luckily all you have to do is find the slope of the line through those two points

jimthompson5910 (jim_thompson5910):

using the slope formula m = (y2 - y1)/(x2 - x1)

OpenStudy (anonymous):

(-3^2 - 7^1) / (-1^2 - 2^1)=5.3^_ ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

i should have stated that the 2's and 1's aren't exponents

jimthompson5910 (jim_thompson5910):

they're just ways to label the two sets of x and y values (so you can tell them apart)

OpenStudy (anonymous):

oh ok

jimthompson5910 (jim_thompson5910):

slope through (-3, 7) and (-1, 2) m = (y2 - y1)/(x2 - x1) m = (2 - 7)/(-1 - (-3)) m = (2 - 7)/(-1 + 3) m = ???

OpenStudy (anonymous):

-2.5

jimthompson5910 (jim_thompson5910):

you got a negative slope (which is correct), so you have a negative correlation it's actually a perfect negative correlation because there are only two points on this scatter plot